How much 5% maple syrup and water mixture must be added to 100% pure maple syrup to produce 5 quarts of syrup that is 10% maple syrup?
Answer by mananth(16946)

percent ---------------- quantity
Maple syrup 5 ---------------- x 0 Quarts
Maple syrup 100----------------5-x Quarts
Total 10.00% ----------------5 Quarts

5x+100 (5-x)=10*5
5x+500-100x=50
5x-100x =50-500
-95 x= -450
/-95
x=4.74 quarts 5.00%

0.26 quarts 100.00%